OpenStudy (anonymous):

Determine the largest acute angle of a right triangle which has legs of length 5 and 6: Determine the largest acute angle of a right triangle which has legs of length 5 and 6: @Mathematics

OpenStudy (anonymous):

(alpha) | \ 6 | \ x = 7.81 |___\ (beta) 5 \[x = \sqrt{5^2+6^2} = 7.81\]\[\sin^{-1} (\frac {5}{7.81}) = \alpha\]\[\cos^{-1} (\frac {6}{7.81}) = \beta\]

OpenStudy (anonymous):

ok let me try to enter this into my homework

OpenStudy (anonymous):

i got 39.80 for both and it didnt work

OpenStudy (radar):

Tan beta =6/5=1.2 \[\tan ^{-1}1.2=50.19 degrees\]

OpenStudy (radar):

The little -1 above tan means an angle whose tangent is 1.2

OpenStudy (anonymous):

hey radar, thanks for coming around to help me again. I entered the answer my home work program rejected it. I dont know why

OpenStudy (radar):

I tried the way mathmind worked it and got 39.8 degrees, but that is the "smallest" acute angle. Subtract that from 90 and you will get the largest acute angle, and the problem is asking for the largest acute angle.

OpenStudy (anonymous):

yeah i see what you're saying. i dont why it isnt accepting the answer

OpenStudy (radar):

The answer is 50.2 rounded off to nearest tenth. (degrees)

OpenStudy (anonymous):

it rejected that also

OpenStudy (radar):

I hate rejection! What are the exact instructions for this problem?

OpenStudy (anonymous):

i copied and paste the exact question into this thread title. there is nothing more Determine the largest acute angle of a right triangle which has legs of length 5 and 6:

OpenStudy (radar):

Well I don't know what to say the two angles must add up to 90 and the 39.8 and 50.2 works and the largest is the 50.2.

OpenStudy (anonymous):

yeah i totally see what you're saying. it makes sense

OpenStudy (anonymous):

can we finish the last prob we were working on before you left. i was never able to find AB

OpenStudy (radar):

How do I get back to it?